OpenStudy (anonymous):

Evaluate. a = 6, b = 4, c = 8, and d = 5 (3a)^2+cd-4b

OpenStudy (anonymous):

348

OpenStudy (anonymous):

a = 6, b = 4, c = 8, and d = 5 (3a)^2+cd-4b a = 6, b = 4, c = 8, and d = 5 (3x6)^2+(8x5)-4(4)= 18^2+40-16= 348